The Multiplier in the Simple Model with No Income Taxes

The simple government purchase multiplier is given by 1/(1 - MPC). Suppose that the U.S. MPC is 0.9. Calculate the government purchase multiplier. If the U.S. government increases its spending by $100 billion to stimulate the economy devastated by the Coronavirus crisis, how much will U.S. GDP increase in the simple model?
Answer #1

Answer

Multiplier =1/(1-MPC)

=1/(1-0.9)

=10

the government purchase multiplier is 10

Change in GDP =change in government spending * multiplier

=100*10

=$1000 billion

the GDP increases by $1000 billion